OpenStudy (anonymous):

The integral of (x^2 - x)/(x^2 + x +1) from 0 to 7

OpenStudy (anonymous):

Do I use long division?

myininaya (myininaya):

yes

OpenStudy (anonymous):

so I have 1-(1/x^2 +x +1); correct?

myininaya (myininaya):

1-(2x+1)/(x^2+x+1

myininaya (myininaya):

Yu got it?

OpenStudy (anonymous):

no....

myininaya (myininaya):

Try your division again X^2/X^2=1 so 1(x^2+X+1)=X^2+X+1 (X^2-X)-(x^2+X+1)=-2X-1=-(2X+1)

OpenStudy (nikvist):

\[\int\limits_{0}^{7}\frac{x^2-x}{x^2+x+1}dx=\int\limits_{0}^{7}\frac{x^2+x+1-2x-1}{x^2+x+1}dx=\int\limits_{0}^{7}\left(1-\frac{2x+1}{x^2+x+1}\right)dx=\] \[=\int\limits_{0}^{7}dx-\int\limits_{0}^{7}\frac{d(x^2+x+1)}{x^2+x+1}=x|^7_0-\ln{(x^2+x+1)}|^7_0=7-\ln{57}\]

OpenStudy (anonymous):

natural log, instead of log....that's where i went wrong Thanks!